OpenStudy (anonymous):

A line has a slope of 2/3 and a y-intercept of −5. Which equations represent the line? Choose all answers that are correct.

OpenStudy (kat_s123):

This is in the history section :P but, in y intercept form the equation is y=mx+b where m = the slop and b equals the y intercept, so the y intercept equation would be y =2/3x-5. Then standard form is in this form Ax+By= c. Soo why is already on the side that it needs to be so now all you have to do is move x! So you'd subtract x from both sides and you'd get -2/3x+y=-5. If you know the forumulas for the difference lines it's super easy :)

OpenStudy (anonymous):

sorry and thx